Fermi National Laboratory
 

Operation Summary for May 24, 2010 - May 26, 2010

 

Monday May 24

The day shift began with Operations monitoring the Tevatron (TeV) store 7834 (223.13E30) and the Recycler stash (58.7E10), with the Antiproton Source (Pbar) stacking (16.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, with beam to the Fermilab Test Beam Facility (FTBF), and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations monitored systems.

During the evening shift, Operations terminated store 7834 at 10:36 PM and then put the TeV into HEP shot setup.

Tuesday May 25

The midnight shift began with the TeV in shot setup, Operations monitoring the stash (5.5E10), with Pbar stacking (39.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 7835 at 12:05 AM. Operations prepared the Pbar rings and PreVault areas for access at 6:54 AM for a scheduled maintenance period.

During the day shift, Pbar was in access nearly all shift. Pbar resumed stacking at 3:47 PM. Operations reported that a MI-40 power supply LCW leak and subsequent MI power supply interlock loop problems held off the MI beam from 1:05 PM to 3 PM.

During the evening shift, Operations terminated store 7835 at 9:41 PM and put the TeV into shot setup and established store 7836 at 10:59 PM with an average initial luminosity of 283.7E30.

Wednesday May 26

The midnight shift began with Operations monitoring store 7836 (202.18E30) and the stash (51.9E10), with Pbar stacking (9.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations monitored systems.

The Plans for Wednesday

The plans for today are to stack, stash, and store. TeV personnel will adjust the D0 proton and antiproton beams interaction position before store 7836 is terminated. The Run Coordinator will allow an access into CDF and D0 while cryo system technicians replace the TeV sector B1 wet engine flywheel. FTBF will have two test beam experiments starting up later this afternoon: T-1006 and T-995.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/26/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory